Use the information below to answer questions 3.15 to 3.16.

If 𝐶̅= R20 billion, I = R30 billion, G = R25 billion, X = R15 billion, Z = R12 billion, c = 0.75 and t =

0.20

3.15 The value of the multiplier is …

[1] 4.0

[2] 3.5

[3] 2.0

[4] 2.5

3.16 The equilibrium level of income is …

[1] 195

[2] 312

[3] 156

[4] 273



1
Expert's answer
2021-07-08T15:20:49-0400

3.15 The value of the multiplier is "m = \frac{1}{1 - 0.75×(1 - 0.2)} = 2.5."

So, the correct answer is [4].

3.16 The equilibrium level of income is:

Y = C + G + I + NX = 20 + 0.75(Y - 0.2Y) + 25 + 30 + (15 - 12) = 0.6Y + 78,

0.4Y = 78,

Y = 195.

So, the correct answer is [1].